Object A travels in a straight line and has 100.0J of kinetic energy. It collides with a stationary object B of mass 3.0kg. After the collision assumed to be elastic, object A has 4.0J of kinetic energy. Determine the speed of object B after the collision.
1
Expert's answer
2020-09-15T10:04:50-0400

In elastic collisions the kinetic energy conserves. This means than energies before and after the collision are equal:


"K_A +K_B = K'_A + K'_B"

where "K_A = 100J" is energy of the object before the collision, "K_B = 0" is energy of the object B before the collision (because it is stationary), "K_A' = 4J" is the energy of the object A after the collision and "K_B'" is the energy of the object B after the collision. Thus, obtain:


"K_B' = K_A + K_B - K_A' = 100J + 0J - 4J = 96J"

On the other hand, by definition, the kinetic energy is:


"K_B' = \dfrac{m_Bv_B'^2}{2}"

where "m_B = 3kg" is the mass of the object B, "v_B'" is the speed of the object B after the collision. Expressing "v_B'", obtain:


"v_B' = \sqrt{\dfrac{2K_B'}{m_B}}=\sqrt{\dfrac{2\cdot 96}{3}}= 8 m/s"

Answer. 8 m/s.